OpenStudy (anonymous):

y – 2x = 3 y = -2x + 3 y = -2x - 3 y = 2x + 3 y = 2x - 3 2x + 2y = 8 y = x + 8 y = x + 4 y = -x + 8 y = -x + 4

OpenStudy (anonymous):

-_- i so need a medal Answers --- 1. y - 2 x = 3 2. y = -2x + 3 3. y = -2x - 3 4. y = 2 x + 3 5. y = 2 x - 3 6. 2x+2y=8 7. y = x + 8 8. y = -x + 4 9. y = -x+8 10. y = x + 4